OpenStudy (anonymous):

If the hypotenuse of a right triangle is 2√3 units long and one of the legs is √3 units long, then how long is the other leg? a) 3 b) 2√ 3 c) √3

OpenStudy (radar):

\[a^{2}+b ^{2}=c ^{2}\]

OpenStudy (radar):

\[(\sqrt{3})^{2}+(b ^{2})=(2\sqrt{3})^{2}\]

OpenStudy (radar):

Solve for b

OpenStudy (anonymous):

so its √3

OpenStudy (radar):

\[b ^{2}=(2\sqrt{3})^{2}-(\sqrt{3})^{2}\]

OpenStudy (anonymous):

3

OpenStudy (radar):

\[b ^{2}=12-3=9\]

OpenStudy (radar):

Exactly, Your answer is correct

OpenStudy (anonymous):

ok thanks